# Device-independent certification of desirable properties with a confidence interval

### v2

@@ III.1 Martingale-based protocol
$\left\langle I(v)\right\rangle:=\tfrac{\left\langle\beta^{ab}_{xy}\right\rangle}{P_{xy}}=\sum_{a,b,x,y}\beta^{ab}_{xy}P(ab|xy)\overset{\mathcal{H}}{\leq}B_{\mathcal{H}}(\vec{\beta})$ (24)
@@ III.2 The prediction-based-ratio (PBR) protocol
Next, notice that $\vec{f}^{\prime}$ typically cannot be cast in the form of Eq. 6. Consequently, we observe empirically that the $R$ obtained by solving Eq. 29 with $\vec{f}^{\prime}$ in place of $\vec{f}_{\text{reg}}$ gives evidently suboptimal performance (see, e.g., Figs. 10, 13 and 15 in Appendix A for some explicit examples). As such, we shall first regularize [37] $\vec{f}^{\prime}$ to some outer approximation of the quantum set $\mathcal{Q}_{\ell}$ by solving Eq. 29 with $\mathcal{S}_{\mathcal{H}}$ replaced by $\mathcal{Q}_{\ell}$. In our work, $\mathcal{Q}_{\ell}$ is the $-\ell$ outer approximation of the quantum set $\mathcal{Q}$ introduced in [17]. However, one may also consider other approximations [67, 22]. Since all these outer approximations admit an SDP characterization, this regularization process is a conic program (see also [37]). The resulting minimizer, which we call the regularized relative frequency $\vec{f}_{\text{reg}}$ is then fed into Eq. 29 to obtain the desired PBR.
@@ IV.1.1 Negativity certification
Figure 2: Negativity certifiable from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$. For the martingale-based protocol and any given $\mathcal{N}_{0}$ among $\bm{N}_{0}=\{0,0.01,\cdots,0.49\}$, we use Eq. 37 in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{N}_{0}$ from $\bm{N}_{0}$, we solve Eq. 38 by considering the same block size and the level-$3$ outer approximation of $\mathcal{Q}$ introduced in [17]. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{N}(\rho)$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{N}_{0}$ in $\bm{N}_{0}$ such that $\mathcal{H}_{\mathcal{N}(\rho)\leq\mathcal{N}_{0}}$ is rejected with a $p$-value bound being less than or equal to $0.01$. Each data point shown in the plot is an average over $30$ such lower bounds, and the error bar (standard deviation) gives an indication of the spread of the certifiable negativity. To avoid cluttering the plots, in each line, we show only a small number of markers.
@@ IV.1.2 Dimension certification via negativity certification
Figure 4: Negativity certifiable from the data observed in a Bell test generating $\vec{P}_{\text{CGLMP}}$. For details on how the plot is generated, see the caption of Fig. 2, but bearing in mind that for the martingale-based protocol, we now use in Eq. 25 the $B_{\mathcal{H}}$ determined from Eq. 40.
@@ IV.3 Fidelity certification
Figure 6: Certifiable fidelity $\mathcal{F}_{\theta=\frac{\pi}{4}}(\rho_{\text{\tiny SWAP}})$ from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$. For the martingale-based protocol and any given $\mathcal{F}_{0}$ among $\bm{F}_{0}=\{0.5,0.51,\cdots,0.99\}$, we use the $B_{\mathcal{H}}$ determined from Eq. 52 in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{F}_{0}$ from $\bm{F}_{0}$, we solve Eq. 51 by considering the same block size and the level-$2$ outer approximation of $\mathcal{Q}$ introduced in [17]. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{F}_{\theta=\frac{\pi}{4}}(\rho_{\text{\tiny SWAP}})$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{F}_{0}$ in $\bm{F}_{0}$ such that $\mathcal{H}_{\mathcal{F}_{\theta}(\rho_{\text{\tiny SWAP}})\leq\mathcal{F}_{0}}$ is rejected with a $p$-value upper bound less than or equal to $0.01$.
@@ IV.4 Properties certification via Bell-value certification
Figure 8: Certifiable Bell-CHSH violation $\mathcal{S}_{\text{\tiny CHSH}}$ from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$. For the martingale-based protocol and any given $\mathcal{S}_{0}$ among $\bm{S}_{0}=\{2+k\Delta\mathcal{S}\}_{k=0}^{49}$, we use Eq. 54b in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{S}_{0}$ from $\bm{S}_{0}$, we solve Eq. 54 by considering the same block size. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{S}_{\text{\tiny CHSH}}$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{S}_{0}$ in $\bm{S}_{0}$ such that $\mathcal{H}_{\mathcal{S}_{\text{\tiny CHSH}}\leq\mathcal{S}_{0}}$ is rejected with a $p$-value bound being less than or equal to $0.01$. A separate calculation shows that if we impose Eq. 9c in addition to Eq. 54b, one may find a visually indistinguishable difference ($<5\times 10^{-4}$) in the certifiable $\mathcal{S}_{\text{\tiny CHSH}}$.
@@ V Discussion
In this paper, we explain in details how the two aforementioned hypothesis-testing protocols can be adapted for DI certification of desirable properties. Specifically, we illustrate how we can use them to perform DI certification of the underlying negativity [65], local Hilbert space dimension [15], entanglement depth [77, 78], and fidelity to some target two-qubit entangled pure state |$\ket{\psi(\theta)}$. In each of these examples, we further demonstrate how the certifiable property (with a confidence of $99\%$) varies with the number of experimental trials involved, see Fig. 2, Fig. 4, Fig. 6. Even though we have focused on certifying desirable properties of quantum states, as explained above, it should be clear that the protocols can also be applied to certify desirable properties of the measurement devices [22, 23, 28], etc.
@@
In the i.i.d. setting, the PBR protocol is known to be asymptotically optimal (in terms of its confidence-gain rate). However, we see from Figs. 2, 4 and 6 that for a relatively small number of trials and with the right choice of the Bell function, the martingale-based protocol performs equally well, if not better. A similar observation was also noted in [102] where the authors therein compare the PBR method with the Chernoff-Hoeffding bound in determining the success probability of Bernoulli trials. In our case, this is not surprising as the PBR method does not presuppose a Bell-like inequality but rather sacrifices some of the data to determine one. Indeed, if we equip the PBR protocol with the optimized Bell-like inequality right from the beginning, its performance is, as expected, no worse than the martingale-based protocol. See Figs. 10, 11, 14, 13, 15 and 16 in Appendix A for some explicit examples.
@@
Meanwhile, we also see from Figs. 3, 5 and 7 that for several cases that we have investigated, one’s intuitive choice of the Bell function for the martingale-based method can lead to a relatively poor confidence-gain rate, and hence impairs its efficiency to produce a good $p$-value bound, see Figs. 14 and 16. For example, even though the titled CHSH Bell inequality of Eq. 19 is known to self-test all entangled two-qubit pure states |$\ket{\psi(\theta)}$, this choice of the Bell function in the martingale-based method leads to a worse performance (for bounding the target-state-fidelity) compared with using the CHSH Bell function, which, in turn, gives a suboptimal performance compared with that derived from the PBR protocol, see Fig. 7. At this point, it is worth reiterating that both protocols do not require the assumption that the experimental trials are i.i.d, even though we have only given, for simplicity, examples with i.i.d. trials.
@@
Several research directions naturally follow from the present work. Firstly, even though our hypothesis-testing-based approaches enable rigorous DI certification with a confidence interval, by virtue of the techniques involved, one can only make a rather weak certification: out of the many experimental trials, we can be sure that at least one involves a setup that exhibits the desired property (say, with $99\%$ confidence). This is evidently far from satisfactory. A preferable certification scheme should allow one to comment on the general or average behavior of all the measured samples, as has been achieved in [101, 60] for self-testing.
@@
Given that self-testing with a high fidelity is technically challenging, it is still of interest to devise a general recipe for certifying the average behavior of other more specific properties (such as entanglement, steerability, etc.), which may already be sufficient for the specific information processing task at hand. However, note that the rejection of a null hypothesis on the average behavior (e.g., average negativity $\overline{\mathcal{N}(\rho)}\leq\mathcal{N}_{0}$) necessarily entails the rejection of the corresponding null hypothesis for all trials (e.g., $\mathcal{N}(\rho)\leq\mathcal{N}_{0}$ in every trial). Thus, we may expect a tradeoff when switching from the current kind of hypothesis testing to that for an average behavior.
@@
Also worth noting that if the i.i.d. assumption is somehow granted, then our protocols also certify the quality of the setup for every single runs, including those that have not been measured. In this case, once a sufficiently small $p$-value bound is obtained, one can stop measuring the rest of the systems and use them, instead, for the information processing tasks of interest. Of course, since the i.i.d. assumption is generally not warranted, a protocol that achieves certification for some fraction of the copies while leaving the rest useful for subsequent tasks will be desirable. This has been considered for one-shot distillable entanglement in [20] and the self-testing fidelity in [60]. Again, a general treatment will be more than welcome (see, e.g. [103]).

### v3

@@ III.1 Martingale-based protocol
$\left\langle I(v)\right\rangle:=\left\langle\tfrac{\beta^{ab}_{xy}}{P_{xy}}\right\rangle=\sum_{a,b,x,y}\beta^{ab}_{xy}P(ab|xy)\overset{\mathcal{H}}{\leq}B_{\mathcal{H}}(\vec{\beta})$ (24)
@@
{mdframed}
Inputs: Property to be certified $\mathcal{P}$, raw data $\mathbf{v}=\{v_{i}=(a_{i},b_{i},x_{i},y_{i})\}_{i=1}^{N_{\text{tot}}}$ acquired in a Bell test, a preset confidence level $\gamma$ (we use $\gamma=99\%$ in all the examples below), a Bell-like inequality, cf. Eq. 24, compatible with the Null Hypothesis $\mathcal{H}_{\neg\mathcal{P}}$ (associated with the negation of $\mathcal{P}$).
Output: A $p$-value bound $\mathfrak{p}$ for the Null Hypothesis $\mathcal{H}_{\neg\mathcal{P}}$ to be compatible with $\mathbf{v}$.
Steps:
  1. 1.
    Compute $\mathfrak{b}_{\pm}$ from Eq. 26 and the average value $\hat{I}(\mathbf{v})=\frac{1}{N_{\text{tot}}}\sum_{i=1}^{N_{\text{tot}}}I(v_{i})=\frac{1}{N_{\text{tot}}}\sum_{i=1}^{N_{\text{tot}}}\tfrac{\beta^{a_{i}b_{i}}_{x_{i}y_{i}}}{P_{x_{i}y_{i}}}$.
  2. 2.
    Compute the required $p$-value bound $\mathfrak{p}$ by substituting the determined $\mathfrak{b}_{\pm}$ and $\hat{I}(\mathbf{v})$ into Eq. 25 with $N$ replaced by $N_{\text{tot}}$.
Figure 1: Pseudocode associated with the martingale-based protocol for DI certification. If the $p$-value bound $\mathfrak{p}<1-\gamma$, we reject the Null Hypothesis $\mathcal{H}_{\neg\mathcal{P}}$ and hence certify the desired property $\mathcal{P}$ with confidence $\gamma$.
@@ III.2 The prediction-based-ratio (PBR) protocol
Next, notice that $\vec{f}^{\prime}$ typically cannot be cast in the form of Eq. 6. Consequently, we observe empirically that the $R$ obtained by solving Eq. 29 with $\vec{f}^{\prime}$ in place of $\vec{f}_{\text{reg}}$ gives evidently suboptimal performance (see, e.g., Figs. 12, 15 and 17 in Appendix A for some explicit examples). As such, we shall first regularize [37] $\vec{f}^{\prime}$ to some outer approximation of the quantum set $\mathcal{Q}_{\ell}$ by solving Eq. 29 with $\mathcal{S}_{\mathcal{H}}$ replaced by $\mathcal{Q}_{\ell}$. In our work, $\mathcal{Q}_{\ell}$ is the $-\ell$ outer approximation of the quantum set $\mathcal{Q}$ introduced in [17]. However, one may also consider other approximations [67, 22]. Since all these outer approximations admit an SDP characterization, this regularization process is a conic program (see also [37]). The resulting minimizer, which we call the regularized relative frequency $\vec{f}_{\text{reg}}$ is then fed into Eq. 29 to obtain the desired PBR.
@@
{mdframed}
Inputs: Property to be certified $\mathcal{P}$, raw data $\mathbf{v}=\{v_{i}=(a_{i},b_{i},x_{i},y_{i})\}_{i=1}^{N_{\text{tot}}}$ acquired in a Bell test, a preset confidence level $\gamma$ (we use $\gamma=99\%$ in all the examples below), and the distribution $\{P_{xy}\}_{x,y}$.
Output: A $p$-value bound $\mathfrak{p}$ for the Null Hypothesis $\mathcal{H}_{\neg\mathcal{P}}$ (associated with the negation of $\mathcal{P}$) to be compatible with $\mathbf{v}$.
Steps:
  1. 1.
    Choose an appropriate block size $N_{\text{blk}}$ (we use $N_{\text{blk}}=500$ in the examples below).
  2. 2.
    Set $k=1$ and $t_{0}=1$.
  3. 3.
    Use the block of data $\mathbf{v}=\{v_{i}=(a_{i},b_{i},x_{i},y_{i})\}_{i=1}^{N^{(k)}_{\text{est}}}$ to compute the relative frequency $\vec{f}$ via Eq. 28, where $N^{(k)}_{\text{est}}=kN_{\text{blk}}$ if the trials are expected to be (near) i.i.d.
  4. 4.
    If $\vec{f}$ contains zero(s), apply Eq. 33 to obtain $\vec{f}\,^{\prime}$ with only nonvanishing entries; else, set $\vec{f}\,^{\prime}=\vec{f}$.
  5. 5.
    Solve the optimization of Eq. 29 with $\vec{f}\,^{\prime}$ playing the role of $\vec{f}_{\text{reg}}$ and $\mathcal{S}_{\mathcal{H}}$ replaced by $\mathcal{Q}_{\ell}$ [in most examples below, we use the level $\ell=3$ approximation of $\mathcal{Q}$ introduced in [17] for $\mathcal{Q}_{\ell}$].
  6. 6.
    Solve the optimization of Eq. 29 using the optimizer from the last step as $\vec{f}_{\text{reg}}$ where $\mathcal{S}_{\mathcal{H}}$ is now the set of correlations compatible with $\mathcal{H}_{\neg\mathcal{P}}$.
  7. 7.
    Use $\vec{f}_{\text{reg}}$ and the resulting optimizer $\vec{P}_{\star}$ in Eq. 30 to determine the prediction-based-ratio for the $k$-th iteration, $\{R_{k}(a,b,x,y)\}_{a,b,x,y}$.
  8. 8.
    Use $R_{k}$ and the next block of data $\mathbf{v}=\{v_{i}=(a_{i},b_{i},x_{i},y_{i})\}_{i=N^{(k)}_{\text{est}}+1}^{N^{(k)}_{\text{est}}+N^{(k)}_{\text{test}}}$ in Eq. 35 to determine the test statistic $t_{k}$ for the $k$-th iteration. Here, $N^{(k)}_{\text{test}}=\min\{N_{\text{blk}},N_{\text{tot}}-N^{(k)}_{\text{est}}\}$.
  9. 9.
    Unless all $v_{i}$ have been used in the computation, increase $k$ by 1.
  10. 10.
    Repeat steps 3 to 9 until all $v_{i}$ have been used in the computation.
  11. 11.
    Compute the required $p$-value bound $\mathfrak{p}$ by using the last $t_{k}$ computed as $t(\mathbf{v})$ in Eq. 32.
Figure 3: Pseudocode associated with the PBR protocol for DI certification. If the $p$-value bound $\mathfrak{p}<1-\gamma$, we reject the Null Hypothesis $\mathcal{H}_{\neg\mathcal{P}}$ and hence certify the desired property $\mathcal{P}$ with confidence $\gamma$.
@@ IV.1.1 Negativity certification
Figure 4: Negativity certifiable from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$, which arises by locally measuring the Bell state |$\ket{\psi_{\text{\tiny MES}}}$ of Eq. 15a. For the martingale-based protocol and any given $\mathcal{N}_{0}$ among $\bm{N}_{0}=\{0,0.01,\cdots,0.49\}$, we use Eq. 37 in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{N}_{0}$ from $\bm{N}_{0}$, we solve Eq. 38 by considering the same block size and the level-$3$ outer approximation of $\mathcal{Q}$ introduced in [17]. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{N}(\rho)$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{N}_{0}$ in $\bm{N}_{0}$ such that $\mathcal{H}_{\mathcal{N}(\rho)\leq\mathcal{N}_{0}}$ is rejected with a $p$-value bound being less than or equal to $0.01$. Each data point shown in the plot is an average over $30$ such lower bounds, and the error bar (standard deviation) gives an indication of the spread of the certifiable negativity. To avoid cluttering the plots, in each line, we show only a small number of markers.
@@ IV.1.2 Dimension certification via negativity certification
Figure 6: Negativity certifiable from the data observed in a Bell test generating $\vec{P}_{\text{CGLMP}}$, which arises by locally measuring the partially entangled state of Eq. 39a. For details on how the plot is generated, see the caption of Fig. 4, but bearing in mind that for the martingale-based protocol, we now use in Eq. 25 the $B_{\mathcal{H}}$ determined from Eq. 40.
@@ IV.3 Fidelity certification
Figure 8: Certifiable fidelity $\mathcal{F}_{\theta=\frac{\pi}{4}}(\rho_{\text{\tiny SWAP}})$ from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$, which arises by locally measuring the Bell state |$\ket{\psi_{\text{\tiny MES}}}$ of Eq. 15a. For the martingale-based protocol and any given $\mathcal{F}_{0}$ among $\bm{F}_{0}=\{0.5,0.51,\cdots,0.99\}$, we use the $B_{\mathcal{H}}$ determined from Eq. 52 in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{F}_{0}$ from $\bm{F}_{0}$, we solve Eq. 51 by considering the same block size and the level-$2$ outer approximation of $\mathcal{Q}$ introduced in [17]. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{F}_{\theta=\frac{\pi}{4}}(\rho_{\text{\tiny SWAP}})$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{F}_{0}$ in $\bm{F}_{0}$ such that $\mathcal{H}_{\mathcal{F}_{\theta}(\rho_{\text{\tiny SWAP}})\leq\mathcal{F}_{0}}$ is rejected with a $p$-value upper bound less than or equal to $0.01$.
@@ IV.4 Properties certification via Bell-value certification
Figure 10: Certifiable Bell-CHSH violation $\mathcal{S}_{\text{\tiny CHSH}}$ from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$, which arises by locally measuring the Bell state |$\ket{\psi_{\text{\tiny MES}}}$ of Eq. 15a. For the martingale-based protocol and any given $\mathcal{S}_{0}$ among $\bm{S}_{0}=\{2+k\Delta\mathcal{S}\}_{k=0}^{49}$, we use Eq. 54b in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{S}_{0}$ from $\bm{S}_{0}$, we solve Eq. 54 by considering the same block size. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{S}_{\text{\tiny CHSH}}$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{S}_{0}$ in $\bm{S}_{0}$ such that $\mathcal{H}_{\mathcal{S}_{\text{\tiny CHSH}}\leq\mathcal{S}_{0}}$ is rejected with a $p$-value bound being less than or equal to $0.01$. A separate calculation shows that if we impose Eq. 9c in addition to Eq. 54b, one may find a visually indistinguishable difference ($<5\times 10^{-4}$) in the certifiable $\mathcal{S}_{\text{\tiny CHSH}}$.
@@ V Discussion
In this paper, we explain in detail how the two aforementioned hypothesis-testing protocols can be adapted for DI certification of desirable properties. Specifically, we illustrate how we can use them to perform DI certification of the underlying negativity [65], local Hilbert space dimension [15], entanglement depth [77, 78], and fidelity to some target two-qubit entangled pure state |$\ket{\psi(\theta)}$. In each of these examples, we further demonstrate how the certifiable property (with a confidence of $99\%$) varies with the number of experimental trials involved, see Fig. 4, Fig. 6, Fig. 8. Even though we have focused on certifying desirable properties of quantum states, as explained above, the protocols can also be applied to certify desirable properties of the measurement devices, such as their measurement incompatibility [22, 23, 26, 28], or their similarity to some target measurements [87] or instruments [27], etc. Note, however, that the usefulness of our protocols relies on the possibility of certifying the desired property from a Bell-inequality-violating correlation. To this end, we remind that determining the complete list of quantum properties certifiable in a device-independent manner remains, to our knowledge, an open problem.
@@
In the i.i.d. setting, the PBR protocol is known to be asymptotically optimal (in terms of its confidence-gain rate). However, we see from Figs. 4, 6 and 8 that for a relatively small number of trials and with the right choice of the Bell function, the martingale-based protocol performs equally well, if not better. A similar observation was also noted in [102] where the authors therein compare the PBR method with the Chernoff-Hoeffding bound in determining the success probability of Bernoulli trials. In our case, this is not surprising as the PBR method does not presuppose a Bell-like inequality but rather sacrifices some of the data to determine one. Indeed, if we equip the PBR protocol with the optimized Bell-like inequality right from the beginning, its performance is, as expected, no worse than the martingale-based protocol. See Figs. 12, 13, 16, 15, 17 and 18 in Appendix A for some explicit examples.
@@
Meanwhile, we also see from Figs. 5, 7 and 9 that for several cases that we have investigated, one’s intuitive choice of the Bell function for the martingale-based method can lead to a relatively poor confidence-gain rate, and hence impairs its efficiency to produce a good $p$-value bound, see Figs. 16 and 18. For example, even though the titled CHSH Bell inequality of Eq. 19 is known to self-test all entangled two-qubit pure states |$\ket{\psi(\theta)}$, this choice of the Bell function in the martingale-based method leads to a worse performance (for bounding the target-state-fidelity) compared with using the CHSH Bell function, which, in turn, gives a suboptimal performance compared with that derived from the PBR protocol, see Fig. 9. At this point, it is worth reiterating that both protocols do not require the assumption that the experimental trials are i.i.d, even though we have only given, for simplicity, examples with i.i.d. trials.
Several research directions naturally follow from the present work. Firstly, there are the scalability questions: (1) how do the number of measurement bases and (2) the number of samples scale with the complexity (say, dimension) of the measured system? The former is again closely related to the general viability of the device-independent certification approach, where our understanding is far from complete. As for the latter, we remark it is indeed one of the goals of the present work to shed light on the sample complexity of our hypothesis-testing-based approaches. In some cases, such as the certification from the GHZ correlation, we see that hundreds of trials suffice, but in some others, several tens of thousands may be required to give a satisfactory level of certification. Still, some general understanding of how the sample size scales depend on the properties to be certified and the confidence level will be surely welcome.
@@
Secondly, for experimental trials expected to deviate significantly from being i.i.d., one should choose a much smaller block size $N_{\text{blk}}$ than the one adopted in our analysis using the PBR protocol. Intuitively, we should choose $N_{\text{blk}}$ so that the trials do not differ significantly within each block of data. In fact, for testing against LHV theories, some guidelines have been provided in [61] on choosing $N_{\text{blk}}$. A similar analysis for other DI certifications is clearly desirable. Next, even though our hypothesis-testing-based approaches enable rigorous DI certification with a confidence interval, by virtue of the techniques involved, one can only make a relatively weak certification: out of the many experimental trials, we can be sure that at least one consists of a setup that exhibits the desired property (say, with $99\%$ confidence). This is evidently far from satisfactory. A preferable certification scheme should allow one to comment on the general or average behavior of all the measured samples, as has been achieved in [101, 60] for self-testing.
@@
Given that self-testing with high fidelity is technically challenging, it is still of interest to devise a general recipe for certifying the average behavior of other more specific properties (such as entanglement, steerability, etc.), which may already be sufficient for the specific information processing task at hand. However, note that the rejection of a null hypothesis on the average behavior (e.g., average negativity $\overline{\mathcal{N}(\rho)}\leq\mathcal{N}_{0}$) necessarily entails the rejection of the corresponding null hypothesis for all trials (e.g., $\mathcal{N}(\rho)\leq\mathcal{N}_{0}$ in every trial). Thus, we may expect a tradeoff when switching from the current kind of hypothesis testing to that for an average behavior.
@@
Also worth noting that if the i.i.d. assumption is somehow granted, then our protocols also certify the quality of the setup for every single run, including those that have not been measured. In this case, once a sufficiently small $p$-value bound is obtained, one can stop measuring the rest of the systems and use them, instead, for the information processing tasks of interest. Of course, since the i.i.d. assumption is generally not warranted, a protocol that achieves certification for some fraction of the copies while leaving the rest useful for subsequent tasks will be desirable. This has been considered for one-shot distillable entanglement in [20] and the self-testing fidelity in [60]. Again, a general treatment will be more than welcome (see, e.g. [103]).